OpenStudy (anonymous):

The second term in a geometric sequence is 50. The fourth term in the same sequence is 112.5. What is the common ratio in this sequence?

OpenStudy (anonymous):

also I would appreciate it if you can give me the answer to this question.

OpenStudy (jhannybean):

geometric sequence: \(a_n = a_1r^{n-1}\)

OpenStudy (anonymous):

ok so whats the final answer?

OpenStudy (jhannybean):

One moment.

OpenStudy (anonymous):

ok

OpenStudy (ohohaye):

what math are you in?

OpenStudy (anonymous):

Algebra 2

OpenStudy (bibby):

lazy lazy lazy you make me sick how about putting some effort into the process of LEARNING?

OpenStudy (anonymous):

please go troll somewhere else nubby Unlike it's really urgent for me to get this answer.

OpenStudy (bibby):

You just want an answer. This isn't trolling it's an unhealthy attitude

OpenStudy (bibby):

Jhan is completely wasting her time on some gutter trash like yourself when you won't even read her explanation

OpenStudy (anonymous):

I need to compare my answer to her and you are seriously not helping by using such degrading comments.

OpenStudy (bibby):

post your answer and show your work. that's how it works on stackexchange, that's how it should work here

OpenStudy (bibby):

You seriously think we haven't heard these excuses 100 time`s before?

OpenStudy (ohohaye):

here is a link to someone else who had the same problem http://openstudy.com/study#/updates/530596b6e4b022471a6500d8 Hope it helps

OpenStudy (anonymous):

Thanks but what I need is the answer so I can check if mine is correct.

OpenStudy (bibby):

what's your answer you shmuck

OpenStudy (jhannybean):

\[a_1r^{2-1}=50~,~ a_1r^{4-1} = 112.5\] \[a_n=a_1r^{n-1}\] \[a_1r^{2-1} =a_1r=50\]\[a_1r^{3-1} = a_1r^2 \]\[a_1r^{4-1} = a_1r^3=112.5\] \[a_1=\frac{50}{r}\]\[a_1=\frac{112.5}{r^3}\]Since it is part of the same sequence, you can solve for \(a_1\) for both of these, and set the ratios equal to eachother and solve for \(r\). \[\dfrac{\dfrac{50}{r}}{\dfrac{112.5}{r^3}}=a_1\]

OpenStudy (jhannybean):

\[a_1: \frac{50}{r}=\frac{r^3}{112.5}\]\[r^4=(50\cdot 112.5)\]\[r=\sqrt[4]{5625}\]\[r=(5625)^{1/4}\]

OpenStudy (jhannybean):

I think if it was part of a different sequence, you would not be able to set the ratios equal to eachother because you would have to infer that \(a_1\) might be different.

OpenStudy (anonymous):

well I did do it that way and I got 1.5 but I don't know if it is correct

OpenStudy (jhannybean):

Can you show how you got 1.5?

OpenStudy (jhannybean):

If you solve for that, \(r\approx 8.66\)

OpenStudy (anonymous):

oh would that be the final answer then?

OpenStudy (jhannybean):

Do you have answeer choices you can check with? I'm curious how you got 1.5.

OpenStudy (anonymous):

I multiplied the 50 and 112.5 then divided and got 75 which is 3/4

OpenStudy (jhannybean):

divided by what?

OpenStudy (anonymous):

nvm thanks for helping

OpenStudy (jhannybean):

Hmm.. ok, np.